OpenStudy (anonymous):

Translate the following situation into an inequality. Do not solve. “Janice and Leonard together would like to save more than $3,500. They have saved $425 saved so far. Janice can save $120 more than twice what Leonard can save each month. How much money will they each need to save to reach their goal

OpenStudy (sayan):

Leonard=x Janice=y=2x+120 Inequality is: y+x+425>=3500 2x+120+x+425>=3500 3x>=3500-545 3x>=2955 x>=985 Leonard=985 Janice=2*985+120=2090